# PULL: PU-Learning-based Accurate Link Prediction

### v1

@@ Abstract.
Given an edge-incomplete graph, how can we accurately find the missing links? The link prediction in edge-incomplete graphs aims to discover the missing relations between entities when their relationships are represented as a graph. Edge-incomplete graphs are prevalent in real-world due to practical limitations, such as not checking all users when adding friends in a social network. Addressing the problem is crucial for various tasks, including recommending friends in social networks and finding references in citation networks. However, previous approaches rely heavily on the given edge-incomplete (observed) graph, making it challenging to consider the missing (unobserved) links during training.
@@
In this paper, we propose PULL (PU-Learning-based Link predictor), an accurate link prediction method based on the positive-unlabeled (PU) learning. PULL treats the observed edges in the training graph as positive examples, and the unconnected node pairs as unlabeled ones. PULL effectively prevents the link predictor from overfitting to the observed graph by proposing latent variables for every edge, and leveraging the expected graph structure with respect to the variables. Extensive experiments on five real-world datasets show that PULL consistently outperforms the baselines for predicting links in edge-incomplete graphs.
link prediction, PU-learning, Markov network, GNN
copyright: acmcopyrightjournalyear: 2024doi: 10.1145/1122445.1122456conference: arXiv preprint; 978-1-4503-XXXX-X/18/06isbn: ;
1. Introduction
@@
Given an edge-incomplete graph, how can we accurately find the missing links among the unconnected node pairs? Edge-incomplete graphs are easily encountered in real-world networks. In social networks, connections between users can be missing since we do not check every user when adding friends. In the context of citation networks, there may be missing citations as we do not review all published papers for citation. The objective of the link prediction in edge-incomplete graphs is to discover the undisclosed relationships between examples when we are provided with a graph that represents the known relationships among them (Liben-Nowell and Kleinberg, 2007). Such scenarios include finding uncited references in citation networks (Shibata et al., 2012; Liu et al., 2019), and recommending new friends in social networks (Wang et al., 2015; Daud et al., 2020a).
@@
The main limitation of previous works (Kipf and Welling, 2016b, 2017; Pan et al., 2018; Zhang and Bai, 2023) for link prediction is that they rely strongly on the given edge-incomplete graph. They presume the edges of the given graph are fully observed ones, and do not consider the unobserved missing links while training. However, this does not always reflect the real-world scenarios where the presence of missing edges is frequently observed. This limits the model’s ability to propagate information through unconnected node pairs, which may potentially form edges, resulting in overfitting of a link predictor to the given edge-incomplete graph. Thus, it is important to consider the uncertainties of the given graph to obtain accurate linking probabilities between nodes.
@@
In this work, we propose PULL (PU-Learning-based Link predictor), an accurate method for link prediction in edge-incomplete graphs. To account for the uncertainties in the given graph structure while training a link predictor $f$, PULL exploits PU (Positive-Unlabeled) learning (see Section 2.2 for details). We treat the observed edges within the graph as positive examples and the unconnected node pairs, which may contain hidden connections, as unlabeled examples. We then construct an expected graph $\mathcal{\bar{G}}$ while proposing latent variables for the unlabeled (unconnected) node pairs to consider the hidden connections among them. This enables us to effectively propagate information through the unconnected edges, improving the prediction accuracy of $f$. Since the estimated linking probabilities of $f$ are the prior knowledge for constructing the expected graph structure $\mathcal{\bar{G}}$, the improved link predictor $f$ enhances the quality of $\mathcal{\bar{G}}$. Thus, PULL employs an iterative learning approach with two-steps to achieve a repeated improvement of the link predictor $f$: a) constructing an expected graph structure $\mathcal{\bar{G}}$ based on the linking probabilities between nodes from the link predictor $f$, and b) training $f$ exploiting the expected graph $\mathcal{\bar{G}}$. Note that the updated $f$ is used to update $\mathcal{\bar{G}}$ in the next iteration. The overall structure of PULL is depicted in Figure 1.
@@
  • Method. We propose PULL, an accurate link prediction method in graphs. PULL effectively overcomes the primary limitation of previous methods, specifically their heavy reliance on the provided graph structure. This is achieved by training a link predictor with an expected graph structure while treating the unconnected edges as unlabeled ones.
@@
  • Theory. We theoretically analyze PULL, including the relationship with the EM algorithm and the time complexity.
@@
  • Experiments. We perform extensive experiments on five real-world datasets, and show that PULL achieves the state-of-the-art link prediction performance.
Figure 1. Overall structure of PULL. Given an edge-incomplete graph $\mathcal{G}_{\mathcal{P}}$ with a set $\mathcal{P}$ of observed edges, PULL first computes the expected graph structure $\mathcal{\bar{G}}$ by proposing latent variables for the edges. Then PULL utilizes $\mathcal{\bar{G}}$ to update the link predictor $f$. The marginal linking probabilities $\hat{y}$ obtained by the updated $f$ are used to compute $\mathcal{\bar{G}}$ in the next iteration.
The rest of this paper is structured as follows. Section 2 introduces related works and the formal definition of the problem. Section 3 proposes PULL along with its theoretical properties. Section 4 shows experimental results, and we conclude the paper at Section 5. The symbols used in this paper is in Table 1.
Table 1. Symbols.
Symbol Description
$\mathcal{E}_{\mathcal{P}}$ of observed edges
$e_{ij}$ Edge between nodes $i$ and $j$
$L(\mathcal{G})$ Corresponding line graph of $\mathcal{G}$
$\mathbf{A}^{\mathcal{G}}$ Corresponding adjacency matrix of $\mathcal{G}=(\mathcal{V},\mathcal{E})$
where $\mathbf{A}^{\mathcal{G}}_{ij}=1$ if $e_{ij}\in\mathcal{E}$
$\mathbf{X}$ Feature matrix for every node in $\mathcal{G}_{\mathcal{P}}$
$f_{\theta}(\cdot,\cdot)$ Link predictor parameterized by $\theta$
$\mathcal{L}(\cdot)$ Objective function that PULL aims to minimize
$\bar{\mathcal{G}}$ Expected graph structure
$\bar{\mathcal{G}}^{\prime}$ Approximated version of $\bar{\mathcal{G}}$
2. Related Works and Problem
We present related works on the link prediction in graph and PU learning. Then we formally define the link prediction problem in edge-incomplete graph.
2.1. Link Prediction in Graphs
@@
Link prediction in graphs has garnered significant attention in recent years, due to its successful application in various domains including social networks (Backstrom and Leskovec, 2011; Wang et al., 2016a; Daud et al., 2020b), recommendation systems (Afoudi et al., 2023; Kurt et al., 2019), and biological networks (Sulaimany et al., 2018; Long et al., 2022). Previous approaches for link prediction are categorized into two groups: embedding-based and autoencoder-based approaches.
Embedding-based approaches strive to create compact representations of nodes within a graph via random walk or propagation. These representations are subsequently employed to estimate the probability of connections between nodes. Deepwalk  (Perozzi et al., 2014) and Node2Vec  (Grover and Leskovec, 2016) create embeddings by simulating random walks on the graph. The concept is to generate embeddings in a way that nodes frequently appearing together in these random walks end up having similar representations. GCN (Kipf and Welling, 2017), LINE (Xu, 2017), GraphSAGE (Hamilton et al., 2017), and GAT (Velickovic et al., 2017) aggregate information from neighboring nodes to learn the embeddings, assuming adjacent nodes are similar. SEAL (Zhang and Chen, 2018) extends the link prediction problem into a subgraph classification problem. However, those methods assume that the edges of the given graph are fully observed. This overfits the node embeddings to the given edge-incomplete graph, degrading the link prediction performance.
@@
The autoencoder-based methods exploit autoencoders to train a link predictor. GAE (Kipf and Welling, 2016a) is an autoencoder-based unsupervised framework for link prediction. VGAE is a variational graph autoencoder, which is a variant of GAE. VGAE explicitly models the uncertainty by introducing a probabilistic layer. ARGA and ARGVA (Pan et al., 2018) exploit adversarial training strategy to improve the performance of GAE and VGAE, respectively. VGNAE (Ahn and Kim, 2021) founds that autoencoder-based methods produce embeddings that converge to zero for isolated nodes, regardless of their input features. They utilize L2-normalization to get better embeddings for these isolated nodes. However, those autoencoder-based methods also have limitations in that they cannot consider the missing edges during training.
2.2. PU Learning
The objective of PU (Positive-Unlabeled) learning is to train a binary classifier that effectively distinguishes positive and negative instances when only positive and unlabeled examples are available. Many algorithms are developed to address the uncertainty introduced by the lack of labeled negative examples. Unbiased risk estimator (URE) (du Plessis et al., 2014) considers the probability that each unlabeled example is a positive instance and adjusts the risk estimate accordingly. Non-negative risk estimator (Kiryo et al., 2017) improves the classification accuracy of URE by preventing the risk of unlabeled instances from going negative. However, those risk-based approaches require the ratio of real positive examples (class prior) among the whole ones in advance, which is not realistic.
@@ 2.2. PU Learning
Many graph-based PU learning approaches have been studied recently (Li et al., 2016; Zhang et al., 2019; Wu et al., 2019). PU-LP (Ma and Zhang, 2017) finds relatively positive examples from the unlabeled ones utilizing the given graph structure, and treats the rest as relatively negative ones. GRAB (Yoo et al., 2021) is the first approach to solve the graph-based PU learning problem without knowing the class prior in advance. However, those graph-based PU learning methods cannot be directly used in the link prediction problem since they aim to classify nodes, not edges, while considering the edges of the given graph as fully observed ones. PU-AUC (Hao, 2021) and Bagging-PU (Gan et al., 2022) proposed PU learning frameworks for link prediction considering the given edges as observed positive examples. However, their link prediction performance is constrained by the propagation of information through the edge-incomplete graph for obtaining node and edge representations.
2.3. Problem Definition
Problem 1 (Link Prediction in Edge-incomplete Graphs).
3. Proposed Method
@@
We propose PULL (PU-Learning-based Link predictor), an accurate method for link prediction in edge-incomplete graphs. PULL effectively exploits the missing links for training the link predictor based on PU-learning approach. We illustrate the entire process of PULL in Figure 1 and Algorithm 1. The main challenges and our approaches are as follows:
  1. (1)
@@
    How can we consider the missing links during training? We treat the given edges as observed positive examples, and the rest as unlabeled ones. We then propagate information through an expected graph structure by proposing latent variables to the unconnected edges. (Section 3.1).
  2. (2)
@@
    How can we effectively model the expected graph structure? Computing the expected graph structure is computationally expensive since there are $2^{\mid\mathcal{E}_{\mathcal{U}}\mid}$ possible graph structures where $\mathcal{E}_{\mathcal{U}}$ is the set of unconnected edges. We effectively compute the expectation of graph structure by carefully designing the probabilities of graphs (Section 3.2).
  3. (3)
@@
    How can we gradually improve the performance of the link predictor? PULL iteratively improves the quality of the expected graph structure, which is the evidence for training the link predictor (Section 3.3).
3.1. Modeling Missing Links
@@
In the problem of link prediction in edge-incomplete graph, we are given a feature matrix $\mathbf{X}$ and an edge-incomplete graph $\mathcal{G}_{\mathcal{P}}$ consisting of two sets of edges, $\mathcal{E}_{\mathcal{P}}$ and $\mathcal{E}_{\mathcal{U}}$. The set $\mathcal{E}_{\mathcal{P}}$ contains observed edges, and $\mathcal{E}_{\mathcal{U}}$ consists of unconnected node pairs; $\mathcal{E}_{\mathcal{P}}\cup\mathcal{E}_{\mathcal{U}}$ is a set of all possible node pairs. Then we aim to find unobserved connected edges among $\mathcal{E}_{\mathcal{U}}$ accurately (see Problem 1 in Section 2.3 for a formal problem definition). Existing link prediction methods treat the given edges of $\mathcal{E}_{\mathcal{P}}$ as fully observed ones, and they propagate information through it to train a link predictor $f$. This overfits $f$ to the edge-incomplete graph, degrading the link prediction performance.
@@
To prevent the overfitting of $f$ to the edge-incomplete graph, PULL models the given graph based on PU-learning approach. Since there are hidden connections in $\mathcal{E}_{\mathcal{U}}$, we treat the unconnected edges in $\mathcal{E}_{\mathcal{U}}$ as unlabeled examples, and the observed edges in $\mathcal{E}_{\mathcal{P}}$ as positive ones. Then we propose a latent variable $z_{ij}\in\{1,0\}$ for every edge $e_{ij}$, indicating whether there is a link between nodes $i$ and $j$ to consider the hidden connections; $z_{ij}=1$ for every $e_{ij}\in\mathcal{E}_{\mathcal{P}}$, but not always $z_{ij}=0$ for $e_{ij}\in\mathcal{E}_{\mathcal{U}}$. We denote the graph $\mathcal{G}_{\mathcal{P}}$ with latent variable $\mathbf{z}=\{z_{ij}\mathrm{\;for\;}e_{ij}\in(\mathcal{E}_{\mathcal{P}}\cup\mathcal{E}_{\mathcal{U}})\}$ as $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$.
@@
A main challenge is that we cannot propagate information through the variablized graph $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$ while training $f$ since every edge $e_{ij}\in\mathcal{E}_{\mathcal{U}}$ of $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$ is probabilistically connected. Instead, PULL exploits the expected graph structure $\bar{\mathcal{G}}$ over the latent variables $\mathbf{z}$. This enables us to train a link predictor $f$ accurately, considering the hidden connections in $\mathcal{E}_{\mathcal{U}}$. Since the link predictor gives prior knowledge for constructing the expected graph, improved $f$ enhances the quality of $\mathcal{\bar{G}}$. Thus, PULL trains the link predictor $f$ by iteratively performing the two steps: a) constructing an expected graph structure $\bar{\mathcal{G}}$ given marginal linking probabilities of trained $f$, and b) updating the link predictor $f$ using $\bar{\mathcal{G}}$, which is used to improve the quality of the expected graph in the next iteration. We use a GCN followed by a sigmoid function as the link predictor $f_{\theta}(i,j)$: $f_{\theta}(i,j)=\mathrm{sigmoid}(h_{i}\cdot h_{j})$ where $h_{i}$ is the hidden representation of node $i$ computed by the GCN embedding function with graph $\bar{\mathcal{G}}$. Other graph-based models can also be used instead of GCN (see Section 4.3).
@@
Input : Edge-incomplete graph $\mathcal{G}_{\mathcal{P}}=(\mathcal{V},\mathcal{E}_{\mathcal{P}})$, feature matrix $\mathbf{X}$, set $\mathcal{E}_{\mathcal{U}}$ of unconnected edges, hyperparameter $r$, and link predictor $f_{\theta}(i,j)$ parameterized by $\theta$
Output : Best parameters $\theta$ of link predictor $f_{\theta}(i,j)$

### v2

# Accurate Link Prediction for Edge-Incomplete Graphs via PU Learning

@@ Abstract
Given an edge-incomplete graph, how can we accurately find its missing links? The problem aims to discover the missing relations between entities when their relationships are represented as a graph. Edge-incomplete graphs are prevalent in real-world due to practical limitations, such as not checking all users when adding friends in a social network. Addressing the problem is crucial for various tasks, including recommending friends in social networks and finding references in citation networks. However, previous approaches rely heavily on the given edge-incomplete (observed) graph, making it challenging to consider the missing (unobserved) links.
@@
In this paper, we propose PULL, an accurate link prediction method based on the positive-unlabeled (PU) learning. PULL treats the observed edges in the training graph as positive examples, and the unconnected node pairs as unlabeled ones. PULL prevents the model from blindly trusting the observed graph by proposing latent variables for unconnected node pairs, and leveraging the expected graph structure with respect to these variables. Extensive experiments on real-world datasets show that PULL consistently outperforms the baselines for predicting links in edge-incomplete graphs.
Abstract
We provide additional information including the problem definition, symbols, lemma, proofs, detailed experimental settings, and further experimental results.
Introduction
@@
Given an edge-incomplete graph, how can we accurately find the missing links among the unconnected node pairs? Edge-incomplete graphs are easily encountered in real-world networks. In social networks, connections between users can be missing since we do not check every user when adding friends. In the context of citation networks, there may be missing citations as we do not review all published papers for citation. In these scenarios, the objective is to find uncited references in citation networks (Shibata, Kajikawa, and Sakata 2012; Liu et al. 2019) or to recommend friends in social networks (Wang et al. 2015; Daud et al. 2020).
@@
The main limitation of previous works (Zhang et al. 2021; Zhu et al. 2021; Chamberlain et al. 2023; Ucar 2023) for link prediction is that they assume the unconnected edges in the given graph as true negative examples. In link prediction, the complete set of true unconnected edges are not given. The given graph contains only a subset of ground-truth edges, while the other node pairs remain unlabeled; i.e., true unconnected edges and the unobserved ground-truth edges are mixed without labels. This misconception limits the model’s ability to propagate information through unconnected node pairs, which may potentially form edges, resulting in over-reliance of a link predictor to the edge-incomplete graph. Thus, it is important to consider the uncertainties of the given graph to obtain accurate linking probabilities.
Figure 1: Main challenge of previous works. They cannot consider the hidden unobserved edges in the given graph. PULL treats the unconnected node pairs as unlabeled examples, and utilizes the expectation of graph structure.
@@
In this work, we propose PULL (PU-Learning-based Link predictor), an accurate link prediction method in edge-incomplete graphs. To account for the uncertainties in the given graph while training a link predictor, PULL exploits PU (Positive-Unlabeled) learning (see Related Works for details). We treat the observed edges in the given graph as positive examples and the unconnected node pairs, which may contain hidden edges, as unlabeled examples. We then construct an expected graph while proposing latent variables for the unlabeled (unconnected) node pairs to consider the hidden edges among them. This enables us to effectively propagate information through the unconnected edges, improving the prediction accuracy. The main challenge of previous works and our approach is depicted in Figure 1.
Figure 2: Overall structure of PULL. Given an edge-incomplete graph $\mathcal{G}_{\mathcal{P}}$ with a set $\mathcal{P}$ of observed edges, PULL first computes the expected graph structure $\mathcal{\bar{G}}$ by proposing latent variables for the edges. Then PULL utilizes $\mathcal{\bar{G}}$ to update the link predictor $f$. The marginal linking probabilities $\hat{y}$ obtained by the updated $f$ are used to compute $\mathcal{\bar{G}}$ in the next iteration.
@@
  • Method. We propose PULL, an accurate link prediction method in graphs. PULL effectively overcomes the primary limitation of previous methods, specifically their heavy reliance on the provided edge-incomplete graph.
@@
  • Theory. We theoretically analyze PULL, including the relationship with the EM algorithm and the complexity.
@@
  • Experiments. Extensive experiments on real-world datasets show that PULL achieves the best performance.
Related Works
Link Prediction in Graphs
@@
Many graph convolutional networks have been proposed to perform various tasks in graph-structured data including graph classification (Yoo, Shim, and Kang 2022), node feature estimation (Yoo et al. 2022a), graph generation (Jung, Park, and Kang 2020), and model compression (Kim, Jung, and Kang 2021). Recently, link prediction has garnered significant attention due to its successful application in various domains including social networks (Daud et al. 2020), recommendation systems (Afoudi, Lazaar, and Hmaidi 2023), and biological networks (Long et al. 2022). Previous link prediction approaches are categorized into two groups: autoencoder-based and embedding-based approaches.
@@
Autoencoder methods use autoencoder structure (Hinton and Salakhutdinov 2006) for the link predictor. GAE and VGAE (2017) are the first autoencoder-based unsupervised framework for link prediction. ARGA and ARGVA (2018) exploit adversarial training strategy to improve the performance of GAE and VGAE, respectively. VGNAE (2021) finds that autoencoder-based methods produce embeddings that converge to zero for isolated nodes. They utilize L2-normalization to get better embeddings for these isolated nodes. MVGAE (2023) is a multi-view representation model which considers both global and local topologies. However, those autoencoder-based methods have limitations in that they cannot consider the missing edges during training.
Embedding-based approaches create compact representations of edges through information aggregation or subgraph extraction. GCN (2017), GraphSAGE (2017), and GAT (2017) aggregate information from neighbors to learn the edge embeddings, assuming adjacent nodes are similar. 1D-GNN (2021) formalizes link prediction as a conditional node classification, incorporating the identity of the source node. Neo-GNN (2022) propagates information through the original graph and concatenates structural features such as the count of common neighbors. SEAL (2018) and BUDDY (2023) extend the link prediction problem to a subgraph classification problem. PS2 (2023) automatically identifies optimal subgraphs for different edges. NESS (2023) utilizes static subgraphs during training and aggregates their representations at test time to learn embeddings in a transductive setting. However, the primary limitation of those methods is that they assume the edges of the given graph are fully observed. This misconception restricts the model’s ability to propagate information between unconnected node pairs that might form edges, leading to over-reliance of the link predictor to the edge-incomplete graph.
Graph-based PU Learning
@@
The objective of PU (Positive-Unlabeled) learning is to train a binary classifier that effectively distinguishes positive and negative instances when only positive and unlabeled examples are available (Kiryo et al. 2017; Zhao et al. 2022). Recently, many graph-based PU learning approaches have been studied (Ma and Zhang 2017; Zhang et al. 2019; Wu et al. 2019). GRAB (2021; 2022b) is the first approach to solve the graph-based PU learning problem without knowing the class prior in advance. PULNS (2021) uses reinforcement learning to design an effective negative sample selector. PU-GNN (2023) introduces distance-aware PU loss to achieve more accurate supervision. However, those methods cannot be directly used in the link prediction problem since they aim to classify nodes, not edges, while considering the edges of the given graph as fully observed ones. PU-AUC (2021) and Bagging-PU (2022) proposed PU learning frameworks for link prediction considering the given edges as observed positive examples. However, their performance is constrained by the propagation of information through the edge-incomplete graph for obtaining node and edge representations.
Proposed Method
@@
We propose PULL, an accurate method for link prediction. We illustrate the overall process of PULL in Figure 2 and Algorithm 1. The main challenges and our approaches are:
  1. C1.
@@
    How can we consider the missing links? We treat the given edges as observed positive examples, and the rest (unconnected edges) as unlabeled ones. We then propagate information through expectation of graph structure by proposing latent variables to the unconnected edges.
  2. C2.
@@
    How can we effectively model the expected graph? Naive computation of the expected graph is intractable since there are $2^{\mid\mathcal{E}_{\mathcal{U}}\mid}$ possible structures where $\mathcal{E}_{\mathcal{U}}$ is the set of unconnected edges. We compute the expectation of graph by carefully designing the probabilities of graphs.
  3. C3.
@@
    How can we gradually improve the performance of the link predictor? PULL iteratively improves the quality of the expected graph structure, which is the evidence for training the link predictor.
Modeling Missing Links (C1)
@@
In a link prediction problem, we are given a feature matrix $\mathbf{X}$ and an edge-incomplete graph $\mathcal{G}_{\mathcal{P}}$ consisting of two sets of edges, $\mathcal{E}_{\mathcal{P}}$ and $\mathcal{E}_{\mathcal{U}}$. The set $\mathcal{E}_{\mathcal{P}}$ contains observed edges, while $\mathcal{E}_{\mathcal{U}}$ consists of unconnected node pairs; $\mathcal{E}_{\mathcal{P}}\cup\mathcal{E}_{\mathcal{U}}$ is a set of all possible node pairs. Then we aim to find unobserved connected edges among $\mathcal{E}_{\mathcal{U}}$ accurately. Existing link prediction methods treat $\mathcal{E}_{\mathcal{U}}$ as true negative examples, which restricts the model’s ability to propagate information through unconnected node pairs that could potentially form edges. This misleads $f$ into fitting the edge-incomplete graph, thereby degrading the prediction performance.
@@
PULL addresses this misconception of previous methods by modeling the given graph based on PU-learning. Since there are hidden connections in $\mathcal{E}_{\mathcal{U}}$, we treat the unconnected edges in $\mathcal{E}_{\mathcal{U}}$ as unlabeled examples, and the observed edges in $\mathcal{E}_{\mathcal{P}}$ as positive ones. Then we propose a latent variable $z_{ij}\in\{1,0\}$ for every edge $e_{ij}$, indicating whether there is a link between nodes $i$ and $j$ to consider the hidden connections; $z_{ij}=1$ for every $e_{ij}\in\mathcal{E}_{\mathcal{P}}$, but not always $z_{ij}=0$ for $e_{ij}\in\mathcal{E}_{\mathcal{U}}$. We denote the graph $\mathcal{G}_{\mathcal{P}}$ with latent variable $\mathbf{z}=\{z_{ij}\mathrm{\;for\;}e_{ij}\in(\mathcal{E}_{\mathcal{P}}\cup\mathcal{E}_{\mathcal{U}})\}$ as $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$.
@@
A main challenge is that we cannot propagate information through the variablized graph $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$ since every edge $e_{ij}\in\mathcal{E}_{\mathcal{U}}$ of $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$ is probabilistically connected. Instead, PULL exploits the expectation $\bar{\mathcal{G}}$ of graph $\mathcal{G}_{\mathcal{P}}(\mathbf{z})$ over the latent variables $\mathbf{z}$. This enables us to train a link predictor $f$ accurately, considering the hidden connections in $\mathcal{E}_{\mathcal{U}}$. Given that the estimated linking probabilities of $f$ provide prior knowledge for constructing the expected graph $\mathcal{\bar{G}}$, improved link predictor $f$ enhances the quality of $\mathcal{\bar{G}}$. Thus, PULL employs an iterative learning approach with two-steps to achieve a repeated improvement of the link predictor: a) constructing an expected graph $\mathcal{\bar{G}}$ based on the predicted linking probabilities from $f$, and b) updating $f$ utilizing $\mathcal{\bar{G}}$. The updated $f$ is used to refine $\mathcal{\bar{G}}$ in the subsequent iteration.
@@
Input : Edge-incomplete graph $\mathcal{G}_{\mathcal{P}}=(\mathcal{V},\mathcal{E}_{\mathcal{P}})$, feature matrix $\mathbf{X}$, set $\mathcal{E}_{\mathcal{U}}$ of unconnected edges, hyperparameter $r$, and link predictor $f_{\theta}$
Output : Best parameters $\theta$ of the link predictor $f_{\theta}$